Equity Lightning Corp. wishes to explore the effect on its cost of capital of the rate at which the company pays taxes. The firm wishes to maintain a capital structure of 30% debt, 10% preferred stock, and 60% common stock. The cost of financing using retained earnings is 14%, the cost of preferred stock financing is 9%, and the before-tax cost of debt financing is 11%. Calculate the weighted average cost of capital (WACC) given the following tax rate assumptions: (1) tax rate=40%, (2) tax rate=35%, (3) tax rate= 25%, describe the effect of tax rate on WACC.

A. (1) 11.28%

(2) 11.45%

(3) 11.78%

As the tax rate decreases, the WACC increases due to the reduced tax shield from the tax-deductible interest on debt

. B. (1) 12.82%

(2) 13.35%

(3) 13.88%

As the tax rate decreases, the WACC increases due to the reduced tax shield from the tax-deductible interest on debt.

C. (1) 11.28%

(2) 11.45%

(3) 11.78%

As the tax rate increases, the WACC increases due to the reduced tax shield from the tax-deductible interest on debt.

D. (1) 11.28%

(2) 11.45%

(3) 11.78%

As the tax rate decreases, the WACC decreases due to the reduced tax expense.
Answer #1

Given Capital Structure was Debt = 30%, Preferred Stock = 10%, Common Stock = 60%

Cost of debt = 11%

Cost of preferred stock = 9%

Cost of retained earnings = 14%

(1) WACC if tax rate was 40%:

Ko = (cost of equity*equity share) + (cost of preferred stock*share of preferred stock) + (cost of debt after tax*share of debt)

Ko = (0.60*0.14) + (0.10*0.09) + (0.30*(0.11)(0.60))

= 0.084 + 0.09 + 0.0198

= 0.1128 or 11.28%

(2) WACC if tax rate was 35%:

Ko = (0.60*0.14) + (0.10*0.09) + (0.30*(0.11)(0.65))

= 0.084 + 0.09 + 0.02145

= 0.11445 or 11.45%

(3) WACC if tax rate was 25%:

Ko = (0.60*0.14) + (0.10*0.09) + (0.30*(0.11)(0.75))

= 0.084 + 0.09 + 0.02475

= 0.1177 or 11.77%

As the tax rate decreases the WACC has increased due to the decreased tax shield from interest on debt

So option A is correct.